Nambe Mills manufactures a line of metal tableware called Nambeware. A piece of Nambeware is produced by sand casting a special alloy of several metals. After casting, the piece goes through a series of steps, including shaping, grinding, buffing, and polishing The top and bottom portions of the data set that Nambe Mills collected for a study of its production schedule are reproduced below; dots indicate that the intervening rows in the data set are not displayed. (Source: DASL [The Data and Story Library]) Piece ID Type Diameter (Inches) Polishing Time (Minutes) 47.65 63.13 58.76 34.88 55.53 Price (Dollars) 144 215 105 69 134 14 10 56 57 58 59 20.85 26.25 21.87 23.88 16.66 24.50 52 62.50 89 21.50 15 Note that the variable Type has five different values, which are coded as 1-bowl, 2 casserole, 3-dish, 4 tray, and 5 = plate Complete the table below to indicate the number of elements, variables, observations, and data values that are included in the complete data set. Data Set Term Elements Variables Observations Measurements How Many? Which variables in the data set are quantitative? Diameter and Price Type, Diameter, Polishing Time, and Price Diameter, Polishing Time, and Price Polishing Time and Price
Answer #1

elements =59 (as 59 ID's on which data is collected)

Variables =4 (type , Dia; polishing time ; price)

Observations =59

measurements =59*4=236

quantitiative variables: Diameter ; polishing time and price

> ***CORRECT ANSSWER*** Quantitative variables - type

Ja'na Wed, Jan 11, 2023 7:19 PM